How much pure acid should be mixed with 9 gallons of a 50% acid solution in order to get an 80% acid solution?
Answer by CharlesG2(834)

How much pure acid should be mixed with 9 gallons of a 50% acid solution in order to get an 80% acid solution?
x gallons of 100% + 9 gallons of 50% = 9+x gallons of 80%
1.00 * x + 0.50 * 9 = 0.8 * (9 + x)
x + 4.5 = 7.2 + 0.8x
0.2x = 2.7
x = 13.5 gallons